Let Z denote the standard normal random variable. If P(0 < Z < b) = 0.437, what is the value of b?
Answer #1

Given: ZNNCoil) & PCOLZL61-0-37 i.e dcbi-door 0.37 & C63-0.5= 0-37 cdco=o.s) normal &cbl=0.87 0 Now using standard table & c1